OpenStudy (anonymous):

What is the magnitude and direction of Vector GH if G(2, -2) and H(7, 6)?

OpenStudy (ghazi):

of what? vector?

OpenStudy (anonymous):

magnitude: 9.43 units; direction 32.01° magnitude: 6.4 units; direction: 57.99° magnitude: 9.43 units: direction: 57.99° magnitude: 6.4 units: direction: 32.01°

OpenStudy (anonymous):

@ghazi yes

OpenStudy (ghazi):

well for magnitude use sqrt (6+2)^2+ (5-2)^2

OpenStudy (anonymous):

which is approx. 9.43

OpenStudy (ghazi):

\[\sqrt{(6+2)^2+(5-2)^2}\]

OpenStudy (ghazi):

yes

OpenStudy (ghazi):

sorry at the place of 7 i've take 5

OpenStudy (anonymous):

\[\sqrt{89} \approx9.43\]

OpenStudy (ghazi):

yes it's right

OpenStudy (ghazi):

and direction = tan inverse (8/5)=57.99

OpenStudy (anonymous):

okay thatwas my problem i was doing 5/8 so i was getting answers that was wrong. lol thanks so much!!!

OpenStudy (ghazi):

:)